Grade 11Physical Chemistry

The vapour of Hg absorbs some electrons accelerated by potential difference of 4.5 volts ASA result light is emitted If full energy of each electron is converted into light emitted by single Hg atom Find wave number of light

Devesh
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
=1.602×10−19×4.5V.QEnergy of an accelerated electron =
⇒7.209×10−19J
This energy is completely converted into light.
hc/λ
=7.209×10−19
6.625×10−34×3.0×108λ
=7.209×10−19
1/λ  = 7.209×10−19
m=wave no =3.63×106